OpenStudy (anonymous):

What is the formula for writing how many combinations I can make with (A, B, C, D)?

OpenStudy (anonymous):

it would be a+b+c=d I think

OpenStudy (anonymous):

so 12 is the answer, is there a formula for finding it?

OpenStudy (anonymous):

n! is the formula

OpenStudy (anonymous):

so u have 4 letters, that means 4!=4x3x2x1=24

OpenStudy (anonymous):

i cant have any repetition though

OpenStudy (anonymous):

The answer is: A B C D AB AC AD BC BD CD ABC ABD BCD CDA ABCD

OpenStudy (anonymous):

well these are all the possible out comes

OpenStudy (anonymous):

the answer says 15..

OpenStudy (anonymous):

then u should use the C notation for combunation

OpenStudy (anonymous):

u have to know how many repletion u have to all possibilities then u would get the right answer

OpenStudy (anonymous):

i can't see in the question u can't have repetition. but let me analyse it for u

OpenStudy (anonymous):

ABCD ABDC ACBD ADBC

OpenStudy (anonymous):

THESE ARE FOR THE As ok

OpenStudy (anonymous):

The question is how many possibilities can I make with (A, B, C, D) AB = BA ABC = CBA

OpenStudy (anonymous):

ABCD = ABDC = ACBD = ADBC all the same

OpenStudy (anonymous):

now let's go for the B

OpenStudy (anonymous):

OK BUT PUT THAT NOTATION IN THE QUESTION

OpenStudy (anonymous):

you have to ask the right question to get the right answer

OpenStudy (anonymous):

the question on the top of this page mean n! is the formula

OpenStudy (anonymous):

ah ok

OpenStudy (anonymous):

well i would use

OpenStudy (anonymous):

\[\frac{ n! }{ (n-r)! }\]

OpenStudy (anonymous):

according to the change in the question excluding repition

OpenStudy (anonymous):

repetition* sorry

OpenStudy (anonymous):

what should r be?

OpenStudy (anonymous):

u have to make sure what the question is telling u because if the question is different then u could have another formula such as

OpenStudy (anonymous):

\[\frac{ n! }{ (n-r)!r! }\]

OpenStudy (anonymous):

look let me give u an example famous in text books

OpenStudy (anonymous):

Ok the question is: How many possibilities can you write with (A, B, C, D) where AB = BA, ABCD = BACD etc.

OpenStudy (anonymous):

yes so is how many ways you can put the letters in order, so u reduce it by subtracting the order which is r

OpenStudy (anonymous):

so what is "r"?

OpenStudy (anonymous):

\[\frac{ 4! }{ (4-r)! }\]

OpenStudy (anonymous):

the answer is 15 it is not possible

OpenStudy (anonymous):

if u have 4 teams of basketball or football, and u want them to play against each other only once which means they do not repeat the game, how many games are played?

OpenStudy (anonymous):

i have no idea! this isnt a data management course, i really want to just know the formula for how to calculate this

OpenStudy (anonymous):

the answer is A B C D AB AB AD BC BD CD ABC ABD BCD CDA ABCD = 15

OpenStudy (anonymous):

what is the formula to calculate?

OpenStudy (anonymous):

i gave u the formula

OpenStudy (anonymous):

but what is "r"?

OpenStudy (amistre64):

r is the size of the group that you are choosing right?

OpenStudy (amistre64):

how many groups of 4 can be made from a set of 6 distinct elements? n C r, where n = 6 and r = 4 6 C 4 since nPr is simpler to calculate for me, i do (nPr)/r! = nCr 6.5.4.3 ------ = 3.5 = 15 is 6C4 4.3.2

OpenStudy (amistre64):

how many groups of 0, then groups of 1, then 2, then 3, then 4 can be made using abcd? 4C0 + 4C1 + 4C2 + 4C3 + 4C4